OpenStudy (anonymous):

Find the x-and the y-intercepts and graph the equation plotting the intercepts. -2x+5y=-15

OpenStudy (anonymous):

x=5n y=2n-3

OpenStudy (anonymous):

variable for y is y=1/5(2x-15)

OpenStudy (anonymous):

do u have a graph to plot the intercepts?

OpenStudy (amistre64):

a seldom used form of an equation is the intercept form:\[ax+by=c\] \[\frac{ax}{c}+\frac{by}{c}=\frac{c}{c}\] \[\frac{x}{c/a}+\frac{y}{c/b}=1\] the intercepts are then read off by the value of the denominator underneath